Calculate the amount of ice which will separate on cooling a solution containing 50g of ethylene glycol in 200g of water to -9.3degree C (kf for water= 1.86K kg/mol)

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
Here, depression in freezing point ∆ T f = 9 . 3 0 C , here mass of glycol = 50 gm, then mass of water W =
delta Tf =Kf*m
m is the molality of the solution
delta Tf is the depresion in freezing point
9.3=1.86*50*1000/w*62
w=161.29 gm
Hence, the mass of ice separated from water = 200 -161.29 =38.71 gm
Rishi Sharma
3 years ago
Hii Sarika,
The solution is as follows
Here, depression in freezing point ∆ T f = 9 . 3 0 C
here mass of glycol = 50 gm,
then mass of water W = delta Tf =Kf*m
m is the molality of the solution delta
Tf is the depresion in freezing point
9.3=1.86*50*1000/w*62
w=161.29 gm
Hence, the mass of ice separated from water = 200 -161.29 =38.71 gm
Thanks
I hope above solution will clear your all doubts.
Please feel free to post and ask as much doubts as possible.
All the best.